OpenStudy (anonymous):

The area of a rectangular piece of land is 280 square meters. if the length of the land was 5 meters less and the width was 1 meter more, the shape of the land would be a square.

OpenStudy (anonymous):

Part A: Write an equation to find the width (x) of the land.

OpenStudy (anonymous):

please :(

OpenStudy (domebotnos):

l - 5 = x + 1 because a square sides are equal. Then solve: l - 5 = x + 1 -1 -1 l - 6 = x

OpenStudy (anonymous):

could you help me with one more

OpenStudy (domebotnos):

Let me see.

OpenStudy (domebotnos):

Post it in another question.

OpenStudy (anonymous):

what is the width of the land in meters

OpenStudy (anonymous):

its the same question

OpenStudy (domebotnos):

solve for y l - 5 = x + 1 + 5 + 5 l = x + 6 Plug it into the other equation. l*x = 280 (x + 6)*x = 280 x^2 + 6x = 280 -280 -280 x^2 + 6x - 280 Then factor for the only reason answer. One will be positive the other negative.

OpenStudy (domebotnos):

Can u do the factoring?

OpenStudy (anonymous):

yes thanyou

OpenStudy (domebotnos):

Medal, plz? :D

OpenStudy (anonymous):

first time how do you give a medal

OpenStudy (anonymous):

Can be solved in one Mathematica statement. Refer to the attachment.